The area of a triangle is 578 the base is 4 times the length of its height. Find the base and and height of the triangle.
Answer by Zucchini(70)

height= h
Base= 4h
The equation for finding the area of a triangle is 1/2 times base times height. So, plug in what you know.
578 = (1/2)(h)(4h)
578 = 1/2 · 4h^2
1156= 4h^2
289 = h^2
17 = h
So,the height is 17 and the base is (4 ·17)= 68
Height: 17
Base: 68
P.s. I'm assuming there were no units that were provided, if there were just put them next to the value of the base and value of the height.